# Riesz Representer Fitting under Bregman Divergence: A Unified Framework for Debiased Machine Learning

### v1

@@ 5.1 RKHS
There exist constants $0<\gamma<2$, $0\leq{\bm{\beta}}\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\gamma}$, where $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{F}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
For details on bracketing entropy, see Appendix D and Definition 2.2 in van de Geer (2000).
@@
Suppose that $g$ is $\mu$-strongly convex and there exists a constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumptions 5.1 and 5.2 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $\alpha_{0}\in{\mathcal{A}}^{\text{RKHS}}$, then we have
@@
The proof is provided in Appendix D, following the approach of Kanamori et al. (2012). The parameter $\gamma$ is determined by the function class to which $f_{0}$ belongs.
@@ 5.2 Neural Networks
Second, we provide an estimation error analysis when we use neural networks for $\mathcal{H}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022). We define Feedforward neural networks (FNNs) as follows:
@@
The proof is provided in Appendix E, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
@@ 5.3 Construction of an Efficient Estimator
Either of the followings holds: (i) the hypothesis classes $\mathcal{H}$} and $\mathcal{M}$} belong to the Donsker class, or (ii) $\widehat{\gamma}$ and $\widehat{\alpha}$ are estimated via cross fitting.
For example, the Donsker condition holds when the bracketing entropy of $\mathcal{H}$} is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if the Donsker condition does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross-fitting, used in debiased machine learning (Chernozhukov et al., 2018). In debiased machine learning, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $m\big(W_{i},\widehat{\gamma}\big)+\widehat{\alpha}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)$, the observations $(X_{i},Y_{i})$ are not used to construct $\widehat{\gamma}$ and $\widehat{\alpha}$. For more details, see Chernozhukov et al. (2018).
@@
Suppose that Assumptions 5.1, and 5.45.5 hold. Then, the AIPW estimator converges in distribution to a normal distribution as
@@ 6 Applications
$g^{\text{SQ}}(\alpha)=(\alpha-1)^{2},$
and minimize the corresponding empirical Bregman objective. By substituting $g^{\text{SQ}}$ into (1) and using $m^{\text{ATE}}(X,f)=f(1,Z)-f(0,Z)$, we obtain, up to an additive constant that does not depend on $\alpha$,
$\text{BD}_{g^{\text{SQ}}}(\alpha)={\mathbb{E}}\Big[\alpha(D,Z)^{2}-2\big(\alpha(1,Z)-\alpha(0,Z)\big)\Big].$
@@
$\widehat{\text{BD}}_{g^{\text{SQ}}}(\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Big(\alpha(D_{i},Z_{i})^{2}-2\big(\alpha(1,Z_{i})-\alpha(0,Z_{i})\big)\Big).$
This coincides with Riesz regression in Chernozhukov et al. (2021) and corresponds to LSIF in density ratio estimation (Kanamori et al., 2009). With appropriate choices of $\mathcal{A}$}, it also recovers nearest neighbor matching-based constructions, as discussed in Kato (2025a).
@@
We next use the UKL divergence loss with $C=1$,
$g^{\text{UKL}}(\alpha)=(|\alpha|-1)\log\left(|\alpha|-1\right)-|\alpha|,$
which is the ATE-specific specialization of Section 3.4. To encode the ATE structure and preserve the sign, we model $\alpha$ through a propensity score model:
@@
and $\phi\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the UKL flavored empirical Bregman objective yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r_{\bm{\beta}}(0,Z_{i})-1}\right)+r_{\bm{\beta}}(0,Z_{i})\right)\Bigg).$
By duality, this KL divergence objective is equivalent to a covariate balancing program:
@@
This specification has the practical advantage that $\phi(Z)$ can be chosen independently of $D$, and it enforces the correct signs and nonnegativity of the magnitude components $r_{\bm{\beta}}(1,Z)$ and $r_{\bm{\beta}}(0,Z)$ by construction.
@@
BP-Riesz regression uses Basu’s power divergence with $C=1$ and $\gamma\in(0,\infty)$:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-1\big)^{1+\gamma}-\big(|\alpha|-1\big)}{\gamma}-|\alpha|.$
Plugging $g^{\text{BP}}$ into (1) and using $m^{\text{ATE}}$ yields the empirical objective, up to an additive constant that does not depend on $\alpha$,
$\widehat{\text{BD}}_{g^{\text{BP}}}(\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Bigg(\frac{\big(|\alpha(D_{i},Z_{i})|-1\big)^{\gamma}-1}{\gamma}+\big(|\alpha(D_{i},Z_{i})|-1\big)^{\gamma}+\big(|\alpha(D_{i},Z_{i})|-1\big)^{1+\gamma}-m^{\text{ATE}}\left(X_{i},\,k\,\operatorname{sign}(\alpha(\cdot))\Big(\big(|\alpha(\cdot)|-1\big)^{\gamma}-1\Big)\right)\Bigg),$
where $k\coloneqq 1+1/\gamma$ and $m^{\text{ATE}}(X_{i},f)=f(1,Z_{i})-f(0,Z_{i})$. We then estimate $\alpha^{\text{ATE}}_{0}$ by
@@
A convenient parametric specification that is consistent with Section 4 models $\alpha$ through a power series link for the inverse propensity components:
@@
$r_{\bm{\beta}}(1,Z)\coloneqq 1+\left(1+\frac{{\bm{\phi}}(Z)^{\top}{\bm{\beta}}}{k}\right)^{1/\gamma},\qquad r_{\bm{\beta}}(0,Z)\coloneqq 1+\left(1-\frac{{\bm{\phi}}(Z)^{\top}{\bm{\beta}}}{k}\right)^{1/\gamma},$
on the domain where the above powers are well defined. This specification interpolates between the squared loss and UKL formulations: $\gamma=1$ recovers SQ-Riesz regression, and the limit $\gamma\to 0$ recovers UKL-Riesz regression, as discussed in Section 3. In applications, BP-Riesz regression can mitigate sensitivity to extreme inverse propensity weights while retaining the covariate balancing behavior implied by the dual characterization.
@@ 6.2 AME Estimation
BP-Riesz regression interpolates between squared loss and KL type criteria. For simplicity, we present the unshifted form with $C=0$:
@@
Once we obtain $\widehat{\alpha}^{\text{AME}}$ and an outcome regression estimator $\widehat{\gamma}$, we plug them into the Neyman orthogonal score in Section 2 to form a Neyman targeted estimator of $\theta_{0}^{\text{AME}}$.

### v2

@@ 6.1 RKHS
There exist constants $0<\tau<2$, $0\leq{\bm{\beta}}\leq 1$, $c_{0}>0$, and $A>0$ such that for all $M\geq 1$, it holds that $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})\leq A\left(\frac{M}{\delta}\right)^{\tau}$, where $H_{B}(\delta,{\mathcal{F}}^{\text{RKHS}}_{M},P_{0})$ is the bracketing entropy with radius $\delta>0$ for the function class ${\mathcal{F}}^{\text{RKHS}}_{M}$ and the distribution $P_{0}$.
For details on bracketing entropy, see Appendix E and Definition 2.2 in van de Geer (2000).
@@
Suppose that $g$ is $\mu$-strongly convex and there exists a constant $C>0$ such that $|g^{\prime\prime}(t)|\leq C\quad\forall t\in{\mathbb{R}}$. Assume also that $\zeta^{-1}(0)$ is finite. Suppose that Assumptions 6.1 and 6.2 hold. Set the regularization parameter $\lambda=\lambda_{n}$ so that $\lim_{n\to\infty}\lambda_{n}=0$ and $\lambda^{-1}_{n}=O(n^{1-\delta})$ ($n\to\infty$). If $\alpha_{0}\in{\mathcal{A}}^{\text{RKHS}}$, then we have
@@
The proof is provided in Appendix E, following the approach of Kanamori et al. (2012). The parameter $\tau$ is determined by the function class to which $f_{0}$ belongs.
@@ 6.2 Neural Networks
Second, we provide an estimation error analysis when we use neural networks for $\mathcal{A}$}. Our analysis is mostly based on Kato & Teshima (2021) and Zheng et al. (2022). We define Feedforward neural networks (FNNs) as follows:
@@
The proof is provided in Appendix F, following the approach of Zheng et al. (2022). This result directly implies the minimax optimality of the proposed method when $f_{0}$ belongs to a Hölder class.
@@ 6.3 Construction of an Efficient Estimator
Either of the followings holds: (i) the hypothesis classes $\mathcal{A}$} and $\mathcal{M}$} belong to the Donsker class, or (ii) $\widehat{\gamma}$ and $\widehat{\alpha}$ are estimated via cross fitting.
For example, the Donsker condition holds when the bracketing entropy of $\mathcal{A}$} is finite. In contrast, it is violated in high-dimensional regression or series regression settings where the model complexity diverges as $n\to\infty$. For neural networks, the assumption holds if both the number of layers and the width are finite. However, if these quantities grow with the sample size, the assumption is no longer valid.
Even if the Donsker condition does not hold, we can still establish asymptotic normality by employing sample splitting (Klaassen, 1987). There are various ways to implement sample splitting, and one of the most well-known is cross fitting, used in debiased machine learning (Chernozhukov et al., 2018). In debiased machine learning, the dataset is split into several folds, and the nuisance parameters are estimated using only a subset of the folds. This ensures that in $m\big(W_{i},\widehat{\gamma}\big)+\widehat{\alpha}(X_{i})\left(Y_{i}-\widehat{\gamma}(X_{i})\right)$, the observations $(X_{i},Y_{i})$ are not used to construct $\widehat{\gamma}$ and $\widehat{\alpha}$. For more details, see Chernozhukov et al. (2018).
@@
Suppose that Assumptions 6.1, and 6.46.5 hold. Then, the AIPW estimator converges in distribution to a normal distribution as
@@ 7 Applications
$g^{\text{SQ}}(\alpha)=\alpha^{2},$
and minimize the corresponding empirical Bregman objective. By substituting $g^{\text{SQ}}$ into (1) and using $m^{\text{ATE}}(X,\gamma)=\gamma(1,Z)-\gamma(0,Z)$, we obtain, up to an additive constant that does not depend on $\alpha$,
$\text{BD}_{g^{\text{SQ}}}(\alpha)={\mathbb{E}}\left[\alpha(D,Z)^{2}-2\big(\alpha(1,Z)-\alpha(0,Z)\big)\right].$
@@
$\widehat{\text{BD}}_{g^{\text{SQ}}}(\alpha)\coloneqq\frac{1}{n}\sum_{i=1}^{n}\left(\alpha(D_{i},Z_{i})^{2}-2\big(\alpha(1,Z_{i})-\alpha(0,Z_{i})\big)\right).$
This coincides with Riesz regression in Chernozhukov et al. (2021) and corresponds to LSIF in density ratio estimation (Kanamori et al., 2009). With appropriate choices of $\mathcal{A}$}, it also recovers nearest neighbor matching-based constructions, as discussed in Kato (2025c).
Remark (Linear link function).
A recommended Riesz representer modeling is
$\alpha_{\bm{\beta}}(X)={\bm{\phi}}(X)^{\top}{\bm{\beta}},$
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the SQ-Riesz regression yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
$\widehat{{\bm{\beta}}}$ $\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}\frac{1}{n}\sum_{i=1}^{n}\Bigg(\Big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\Big)^{2}-2\Big({\bm{\phi}}(1,Z)^{\top}-{\bm{\phi}}(0,Z)^{\top}\Big){\bm{\beta}}\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, SQ-Riesz regression is equivalent to the following covariate balancing problem:
$\min_{\alpha\in{\mathbb{R}}^{n}}\$$\sum_{i=1}^{n}\alpha_{i}^{2}$
s.t. $\ \ \left|\sum_{i=1}^{n}\alpha_{i}\phi_{j}(X_{i})+\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\right|\leq\lambda\ \ \ \text{for}\ j=1,2,\dots,p,$
$\widehat{w}_{i}=\begin{cases}\widehat{\alpha}(1,Z_{i})&\text{if }D_{i}=1,\\ -\widehat{\alpha}(0,Z_{i})&\text{if }D_{i}=0.\end{cases}.$
@@
Consider Riesz representer models $\alpha$ such that $\alpha(1,x)>1$ and $\alpha(0,x)<-1$ for all $x$. We next use the UKL divergence loss with $C=1$,
$g^{\text{UKL}}(\alpha)=(|\alpha|-1)\log\left(|\alpha|-1\right)-|\alpha|.$
By substituting $g^{\text{UKL}}$ into (1) and using $m^{\text{ATE}}(X,\gamma)=\gamma(1,Z)-\gamma(0,Z)$, we obtain
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)$ $\coloneqq\mathbb{E}\Big[\log\left(|\alpha(X)|-1\right)+|\alpha(X)|$
$\ \ \ \ \ -\Big(\operatorname{sign}\big(\alpha(1,Z)\big)\log\left(|\alpha(1,Z)|-1\right)-\operatorname{sign}\big(\alpha(0,Z)\big)\log\left(|\alpha(0,Z)|-1\right)\Big)\Big].$
Thus, UKL-Riesz regression estimates $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{UKL}}}(\alpha)+\lambda J(\alpha),$
where
$\text{BD}_{g^{\text{UKL}}}\big(\alpha\big)$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\log\left(|\alpha(X_{i})|-1\right)+|\alpha(X_{i})|$
$\ \ \ \ \ -\Big(\operatorname{sign}\big(\alpha(1,Z_{i})\big)\log\left(|\alpha(1,Z_{i})|-1\right)-\operatorname{sign}\big(\alpha(0,Z_{i})\big)\log\left(|\alpha(0,Z_{i})|-1\right)\Big)\Big).$
This coincides with the tailored loss minimization with $\alpha=\beta=-1$ in Zhao (2019) and corresponds to KLIEP in density ratio estimation (Sugiyama et al., 2008).
Remark (Log link function).
A recommended Riesz representer modeling is
$\alpha_{\bm{\beta}}(X)=\mathbbm{1}[D=1]\Big(1+\exp\big(-{\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big)-\mathbbm{1}[D=0]\Big(1+\exp\big({\bm{\phi}}(X)^{\top}{\bm{\beta}}\big)\Big),$
where ${\bm{\phi}}\colon{\mathcal{X}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the UKL-Riesz regression yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
$\widehat{{\bm{\beta}}}$ $\coloneqq\operatorname*{arg\,min}_{{\bm{\beta}}}\frac{1}{n}\sum_{i=1}^{n}\Bigg(\mathbbm{1}[D_{i}=1]\left(-{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}+1+\exp\big(-{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}\big)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathbbm{1}[D_{i}=0]\left({\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}+1+\exp\big({\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}\big)\right)$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\Big({\bm{\phi}}(1,Z)^{\top}-{\bm{\phi}}(0,Z)^{\top}\Big){\bm{\beta}}\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, UKL-Riesz regression is equivalent to the following covariate balancing problem:
$\min_{{\bm{w}}\in(1,\infty)^{n}}\$$\sum_{i=1}^{n}(w_{i}-1)\log(w_{i}-1)$
s.t. $\ \ \left|\sum_{i=1}^{n}\Big(\mathbbm{1}[D_{i}=1]w_{i}\phi_{j}(X_{i})-\mathbbm{1}[D_{i}=0]w_{i}\phi_{j}(X_{i})\Big)-\Big(\phi_{j}(1,Z_{i})-\phi_{j}(0,Z_{i})\Big)\right|\leq\lambda$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \text{for}\ j=1,2,\dots,p,$
$\widehat{w}_{i}=\begin{cases}\widehat{\alpha}(1,Z_{i})&\text{if }D_{i}=1,\\ -\widehat{\alpha}(0,Z_{i})&\text{if }D_{i}=0.\end{cases}.$
This modeling enforces the correct signs and nonnegativity of the Riesz representer.
Remark (Propensity score modeling).
We can interpret that the Riesz representer model is based on a propensity score model:
@@
and ${\bm{\phi}}\colon{\mathcal{Z}}\to{\mathbb{R}}^{p}$ is a basis function. Under this model, minimizing the UKL flavored empirical Bregman objective yields an estimator that satisfies an automatic covariate balancing property, as discussed in Section 4 and Zhao (2019).
@@
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\mathbbm{1}[D_{i}=0]\left(-\log\left(\frac{1}{r_{\bm{\beta}}(0,Z_{i})-1}\right)+r_{\bm{\beta}}(0,Z_{i})\right)\Bigg)+\frac{1}{a}\lambda\|{\bm{\beta}}\|^{a}_{a}.$
By duality, if $a=1$, this KL divergence objective is equivalent to a covariate balancing program:
@@
This specification has the practical advantage that $\phi(Z)$ can be chosen independently of $D$.
@@
BP-Riesz regression uses Basu’s power divergence with $C=1$ and $\delta\in(0,\infty)$:
$g^{\text{BP}}(\alpha)\coloneqq\frac{\big(|\alpha|-1\big)^{1+\delta}-\big(|\alpha|-1\big)}{\delta}-|\alpha|.$
Plugging $g^{\text{BP}}$ into (1) and using $m^{\text{ATE}}$ yields the empirical objective
$\widehat{\text{BD}}_{g^{\text{BP}}}(\alpha)$ $\coloneqq\frac{1}{n}\sum_{i=1}^{n}\Bigg(\frac{\big(|\alpha(D_{i},Z_{i})|-1\big)^{\delta}-1}{\delta}+|\alpha(D_{i},Z_{i})|\big(|\alpha(D_{i},Z_{i})|-1\big)^{\delta}$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\upsilon\Big(\big(|\alpha(1,X)|-1\big)^{\delta}-\big(|\alpha(0,X)|-1\big)^{\delta}\Big)\Bigg),$
where $\upsilon\coloneqq 1+1/\delta$. We then estimate $\alpha^{\text{ATE}}_{0}$ by
@@
Remark (Power link function).
A convenient parametric specification that is consistent with Section 4 models $\alpha$ through a power link function for the inverse propensity components:
@@
$r_{\bm{\beta}}(1,Z)\coloneqq 1+\left(1+\frac{{\bm{\phi}}(1,Z)^{\top}{\bm{\beta}}}{\upsilon}\right)^{1/\delta},\qquad r_{\bm{\beta}}(0,Z)\coloneqq 1+\left(1-\frac{{\bm{\phi}}(0,Z)^{\top}{\bm{\beta}}}{\upsilon}\right)^{1/\delta},$
on the domain where the above powers are well defined. This specification interpolates between the squared distance and UKL divergence: $\delta=1$ recovers SQ-Riesz regression, and the limit $\delta\to 0$ recovers UKL-Riesz regression, as discussed in Section 3. In applications, BP-Riesz regression can mitigate sensitivity to extreme inverse propensity weights while retaining the covariate balancing behavior implied by the dual characterization.
@@
Consider Riesz representer models $\alpha$ such that $\alpha(1,x)>1$ and $\alpha(0,x)<-1$ for all $x$. BKL-Riesz regression uses BKL divergence with $C=1$:
$g^{\text{BKL}}(\alpha)\coloneqq(|\alpha|-1)\log\big(|\alpha|-1\big)-(|\alpha|+1)\log(|\alpha|+1).$
By plugging $g^{\text{BKL}}$ into (1) and using $m^{\text{ATE}}$, we have the following empirical objective function:
$\widehat{\text{BD}}_{g^{\text{BKL}}}(\alpha)$ $\coloneqq\frac{1}{n}\sum^{n}_{i=1}\Big(\log\left(\frac{|\alpha(X_{i})|-1}{|\alpha(X_{i})|+1}\right)$
$\ \ \ \ \ \ -\Big(\log\left(\frac{\alpha(1,Z_{i})-1}{\alpha(1,Z_{i})+1}\right)+\log\left(\frac{-\alpha(0,Z_{i})-1}{-\alpha(0,Z_{i})+1}\right)\Big)\Big).$
We then estimate $\alpha^{\text{ATE}}_{0}$ by
$\widehat{\alpha}\coloneqq\operatorname*{arg\,min}_{\alpha\in{\mathcal{A}}}\widehat{\text{BD}}_{g^{\text{BKL}}}(\alpha)+\lambda J(\alpha).$
Remark (MLE of the propensity score).
@@ 7.2 AME Estimation
BP-Riesz regression interpolates between squared distance and UKL divergence. For simplicity, we present the unshifted form with $C=0$:
@@
Once we obtain $\widehat{\alpha}^{\text{AME}}$ and an outcome regression estimator $\widehat{\gamma}$, we plug them into the Neyman orthogonal score in Section 2 to form an estimator of $\theta_{0}^{\text{AME}}$.